OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 1) and (4, 13)?

OpenStudy (amistre64):

subtract the point for slope and then plug and play

OpenStudy (amistre64):

(–2, 1) -(4, 13) ------- -6 , -12; slope = y/x; 12/6 = 2

OpenStudy (amistre64):

y-Py = m(x-Px) ; where Px and Py are any point given

OpenStudy (anonymous):

and even after you call him a name...he still gives you the answer....u owe him an apology

OpenStudy (amistre64):

no, not me lol

OpenStudy (amistre64):

the guy i deleted

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

oh...missed that

OpenStudy (amistre64):

and maybe that suspension will hold for a few hours